OpenStudy (anonymous):

Real analysis help. What is an interior point of a set written in logical form using logical connectives?

OpenStudy (anonymous):

here is the definition in words: A point p is an interior point of E if there is a neighborhood N of p such that N is subset of E.

OpenStudy (anonymous):

anyone? :(

OpenStudy (anonymous):

@satellite73 please help

OpenStudy (anonymous):

@Loser66 can you help?

OpenStudy (anonymous):

@ganeshie8 please?

OpenStudy (anonymous):

\[\exists r>0\ such\ that\ (something \ here?)\]

OpenStudy (anonymous):

@zephyr141 ??

OpenStudy (anonymous):

what does "logical connectives" mean?

OpenStudy (anonymous):

"and","or", "not" you those. And quantifiers.

OpenStudy (anonymous):

@myininaya can you help?

OpenStudy (anonymous):

\[\exists \epsilon >0\] such that \[N_{\epsilon}(p)=\{q\in X|d(p, q)<\epsilon\}\subset E\]

OpenStudy (loser66):

what is the set?

OpenStudy (anonymous):

metric space. and E is in that metric

OpenStudy (loser66):

I give you formal example. S ={(x,y, z) 0<x<1 , y^2+z^2<=1} find Int S? Int S= {(x,y,z)| 0<x<1 , y^2+z^2<1} If \(p =(x_0,y_0,z_0)\in Int s\) there exists a ball \(B(p,\varepsilon) \leq Int S\) for \(\varepsilon =min\{x_0, 1-x_0, 1-\sqrt{y_0^2+z_0^2}\}\) this set is open. bd(S) ={(x,y,z) x =0 or x=1 or y^2+z^2 =1} For any \(p=(x_0,y_0,z_0)\) in this set and \(\forall \varepsilon >0\), B(p,\(\varepsilon\)) contains points in S and points in \(S^c\) therefore, S\bd S = Int S and Ext S ={(x,y,z) | s<0 or x >1 or y^2+z^2 >1}

OpenStudy (loser66):

my computer is crazy, it gets many virus; have to log off now.

OpenStudy (anonymous):

thank you both :)